# Solving Copyright Infringement on Short Video Platforms: Novel Datasets and an Audio Restoration Deep Learning Pipeline

### v2

@@ 1 Introduction
Figure 1: Comparison of generic videos and arbitrarily BGM-added videos
@@ 2.3 Video-Music Retrieval Approaches
To address these challenges, cross-modal retrieval methods emerged as a powerful alternative. Unlike traditional multimodal approaches, which process audiovisual features independently and demand significant computational resources, cross-modal methods embed video and audio data into a shared latent space. This reduces the complexity of feature extraction and alignment while maintaining high performance Gu et al. (2023); Li and Kumar (2019). For example, UT-CMVMR Mao et al. (2024) employs Optical Flow to capture temporal video dynamics and Rhythm Quantization to analyze audio patterns, enabling efficient video-music alignment even in dynamic environments. These methods are well-suited for short video platforms where metadata is sparse.
@@ 3.2.1 Preprocessing
3.3 Pipeline Performance Evaluation Dataset (OSVAR-160)
@@ 4 Method
4.0.1 Initial Audio Separation: Background Music and Speech Extraction
@@
(1) Background Music Separator: The Background Music Separator utilizes the U-Net–based Demucs model Défossez et al. (2019) to decompose audio into its constituent tracks, such as vocals, bass, and drums. Demucs operates in the time domain, leveraging recurrent layers to capture long-term dependencies Ronneberger et al. (2015), which ensures high-quality separation. The separated BGM components are then consolidated into a single “Mixed-Music” track using Pydub, creating a unified input for subsequent modules.
@@
(2) Speech Extractor: The Speech Extractor applies the Pyannote Audio VAD Bredin et al. (2020) to the vocal output generated by Demucs, accurately detecting and isolating speech. By leveraging recurrent neural networks (RNNs) and temporal convolutional networks (TCNs) Soo Kim and Reiter (2017), the module processes overlapping audio frames with contextual embeddings, ensuring robust extraction of dialogue even in noisy or complex audio environments. This precise speech isolation is crucial for maintaining the integrity of spoken content during BGM removal.
@@
4.0.2 Mixed-Music Separator
@@
4.0.3 Video-Music Matching
@@ 5.1.1 Mixed-Music Separator
[…]
@@ 6 Results
6.0.1 Mixed-Music Separator
@@
6.0.2 Video-Music Matching
@@ 6.0.3 Pipeline Performance
When comparing arbitrary BGM separation to the dialogue-only audio scenario (which achieves an average SDR of 131.24 dB and SI-SDR of 14.00 dB), ours demonstrates slightly better performance. This showcases the model’s ability not only to effectively remove the BGM but also to reconstruct the OST in a manner that closely matches the original audio, significantly preserving essential spectral and temporal details.
These results validate our effectiveness in maintaining both high-quality and structural similarity across diverse audio scenarios. By comparing the restored outputs to the original reference audio, the evaluation ensures that SDR and SI-SDR metrics reliably reflect the model’s capacity to accurately reconstruct audio components.

### v3

@@ 1 Introduction
Figure 1: Comparison of generic and arbitrarily BGM-added videos
@@ 2.3 Video-Music Retrieval Approaches
To address these challenges, cross-modal retrieval methods emerged as a powerful alternative. Unlike traditional approaches, which process audiovisual features independently and demand significant computational resources, cross-modal methods embed video and audio data into a shared latent space. This reduces the complexity of feature extraction and alignment while maintaining high performance Gu et al. (2023); Li and Kumar (2019). For example, UT-CMVMR Mao et al. (2024) employs Optical Flow to capture temporal video dynamics and Rhythm Quantization to analyze audio patterns, enabling efficient video-music alignment even in dynamic environments. These methods are well-suited for short video platforms where metadata is sparse.
@@ 3.2.1 Preprocessing
3.3 Pipeline Performance Evaluation Dataset
@@ 4 Method
4.1 Initial Audio Separation: Background Music and Speech Extraction
@@
(1) Background Music Separator:
The Background Music Separator utilizes the U-Net–based Demucs model Défossez et al. (2019) to decompose audio into its constituent tracks, such as vocals, bass, and drums. Demucs operates in the time domain, leveraging recurrent layers to capture long-term dependencies Ronneberger et al. (2015), which ensures high-quality separation. The separated BGM components are then consolidated into a single “Mixed-Music” track using Pydub, creating a unified input for subsequent modules.
@@
(2) Speech Extractor:
The Speech Extractor applies the Pyannote Audio VAD Bredin et al. (2020) to the vocal output generated by Demucs, accurately detecting and isolating speech. By leveraging recurrent neural networks (RNNs) and temporal convolutional networks (TCNs) Kim and Reiter (2017), the module processes overlapping audio frames with contextual embeddings, ensuring robust extraction of dialogue even in noisy or complex audio environments. This precise speech isolation is crucial for maintaining the integrity of spoken content during BGM removal.
@@
4.2 Mixed-Music Separator
@@
4.3 Video-Music Matching
@@ 5.1.1 Mixed-Music Separator
$\text{SI-SDR}=10\cdot\log_{10}\left(\frac{\|\alpha\cdot S_{t}\|^{2}}{\|\alpha\cdot S_{e}\|^{2}}\right)$ (2)
@@ 6 Results
6.1 Mixed-Music Separator
@@
6.2 Video-Music Matching
@@ 6.3 Pipeline Performance
When comparing arbitrary BGM separation to the dialogue-only audio scenario (which achieves an average SDR of 131.24 dB and SI-SDR of 14.00 dB), ours demonstrates slightly better performance. This showcases the model’s ability not only to effectively remove the BGM but also to reconstruct the OST in a manner that closely matches the original audio, significantly preserving essential spectral and temporal details. These results validate our effectiveness in maintaining both high-quality and structural similarity across diverse audio scenarios. By comparing the restored outputs to the original reference audio, the evaluation ensures that SDR and SI-SDR metrics reliably reflect the model’s capacity to accurately reconstruct audio components.